The information presented here represents selected data from the December 31, 2016, balance sheets and income statements for the year then ended for three firms. Required Calculate the missing amounts for each firm. Firm A Firm B Firm C Total assets, 12/31/16 Total liabilities, 12/31/16 Paid-in capital, 12/31/16 Retained earnings, 12/31/16 Net income for 2016 Dividends declared and paid during 2016 Retained eanings, 1/1/16 $217,500 260,000 102,500 70,000 40,000 27,500 29,500 93,000 55,000 34,000 6,000 25,000 40,500 14,000 62,000
Answer #1
Firm A Firm B Firm C
Total assets, 12/31/16 $ 1,20,500 $     2,17,500 $ 2,60,000
Total Liabilities, 12/31/16 $     40,000 $         95,000 $ 1,02,500
Paid - in - capital, 12/31/16 $     27,500 $         29,500 $     70,000
Retained Earnings, 12/31/16 $     53,000 $         93,000 $     87,500
Net income for 2016 $     34,000 $         55,000 $     40,500
Dividends paid and declared during 2016 $       6,000 $         24,000 $     14,000
Retained earnings, 1/1/16 $     25,000 $         62,000 $     61,000
Workings:
Firm A:
Retained Earnings, 12/31/16 = Retained earnings, 1/1/16 + Net income for 2016 - Dividends paid and declared during 2016
= $         25,000 + $         34,000 - $              6,000
= $         53,000
Total assets, 12/31/16 = Total Liabilities, 12/31/16 + Paid - in - capital, 12/31/16 + Retained Earnings, 12/31/16
= $         40,000 + $         27,500 + $            53,000
= $     1,20,500
Firm B:
Retained Earnings, 12/31/16 = Retained earnings, 1/1/16 + Net income for 2016 - Dividends paid and declared during 2016
$                                                                    93,000 = $         62,000 + $         55,000 - Dividends paid and declared during 2016
Dividends paid and declared during 2016 = $         24,000
Total assets, 12/31/16 = Total Liabilities, 12/31/16 + Paid - in - capital, 12/31/16 + Retained Earnings, 12/31/16
$                                                                2,17,500 = Total Liabilities, 12/31/16 + $         29,500 + $            93,000
Total Liabilities, 12/31/16 = $         95,000
Firm C:
Total assets, 12/31/16 = Total Liabilities, 12/31/16 + Paid - in - capital, 12/31/16 + Retained Earnings, 12/31/16
$                                                                2,60,000 = $     1,02,500 + $         70,000 + Retained Earnings, 12/31/16
Retained Earnings, 12/31/16 = $         87,500
Retained Earnings, 12/31/16 = Retained earnings, 1/1/16 + Net income for 2016 - Dividends paid and declared during 2016
$                                                                    87,500 = Retained earnings, 1/1/16 + $         40,500 - $            14,000
Retained earnings, 1/1/16 = $         61,000